OpenStudy (anonymous):

Find the range of f(x)=3g-5 for the domain (-1.5,2,4). {8, 10, 1, 1} (9.5,-1,-7) (-8,10,-1,-1) (-9.5, 1,7)

OpenStudy (anonymous):

@LynFran

OpenStudy (lynfran):

ok range = y- values domain = x-values

OpenStudy (anonymous):

ok. Thats what I get confused on is finding the domain

OpenStudy (anonymous):

What is it?

OpenStudy (lynfran):

3(-1.5)-5=-9.5 3(2)-5=1 3(4)-5=7